OpenStudy (anonymous):

find all solutions in the interval [0,2pi) Cos 2x + ✔️2 sin X=1

OpenStudy (anonymous):

\[\cos 2x=1-2\sin ^2x\] \[1-2\sin ^2x+2\sin x=1,-2\sin x \left( \sin x-1 \right)=0\] either sinx=0 or sin x=1 x=?